Timon and pumba are brothers , pumba is 3 times older as timon in 5 years pumba will be twice timons age how old are they now

Answer by Cromlix(4381)

Pumba age = 3*Timon age
P = 3T
In 5 years:
Pumba will be 3T + 5
Timon will be T + 5
And Pumba will be twice Timon's age.
3T + 5 = 2(T + 5)
3T + 5 = 2T + 10
3T - 2T = 10 - 5
T = 5
Pumba = 20 years old
Timon = 10 years old.
Hope this helps.
:-)